# What Makes a Good Speech Tokenizer for LLM-Centric Speech Generation? A Systematic Study

### v2

@@ Abstract
Speech-language models (SLMs) offer a promising path toward unifying speech and text understanding and generation. However, challenges remain in achieving effective cross-modal alignment and high-quality speech generation. In this work, we systematically investigate the role of speech tokenizer designs in LLM-centric SLMs, augmented by speech heads and speaker modeling. We compare coupled, semi-decoupled, and fully decoupled speech tokenizers under a fair SLM framework and find that decoupled tokenization significantly improves alignment and synthesis quality. To address the information density mismatch between speech and text, we introduce multi-token prediction (MTP) into SLMs, enabling each hidden state to decode multiple speech tokens. This leads to up to 12 faster decoding and a substantial drop in word error rate (from 6.07 to 3.01). Furthermore, we propose a speaker-aware generation paradigm and introduce RoleTriviaQA, a large-scale role-playing knowledge QA benchmark with diverse speaker identities. Experiments demonstrate that our methods enhance both knowledge understanding and speaker consistency.
Code
https://github.com/cnxupupup/SLM-Decoupled-MTP
Demo
https://cnxupupup.github.io/SLM-Decoupled-MTP-Demo
@@ 1 Introduction
In recent years, large language models (LLMs) have achieved remarkable progress in text understanding and generation (Grattafiori et al. 2024; Qwen et al. 2025), fueling rapid advancements in multimodal models such as vision-language models (Liu et al. 2023), vision-language-action models (Kim et al. 2024), and speech-language models (SLMs) (Chu et al. 2023; Li et al. 2025). Among them, SLMs are particularly valuable for applications like natural and fluent human-computer dialogue (Défossez et al. 2024; Xu et al. 2025) and personalized speech generation (Du et al. 2024a, c). The core challenge lies in generalizing the language understanding and generation capabilities of LLMs to incorporate and integrate speech understanding and generation.
Current approaches for connecting text and speech in SLMs can be broadly categorized into two perspectives. Speech-centric models (Du et al. 2024a, c; Anastassiou et al. 2024) are built upon mature speech generation architectures, with LLMs introduced as conditional context. However, while this approach achieves high-quality speech synthesis, it struggles to fully leverage LLMs’ rich world knowledge and powerful capabilities. LLM-centric models (Défossez et al. 2024; Zhang et al. 2023; Zhan et al. 2024) take LLMs as the backbone and extend them with speech interaction. In theory, it fully utilizes LLMs’ strengths (e.g., intent understanding, instruction following, planning, and decision-making). However, the quality of speech generation often suffers due to challenges in cross-modal alignment. This work focuses on the LLM-centric SLMs, aiming to explore how core components affect cross-modal alignment and the quality of generated speech.
@@
We observe a mismatch in information density between speech and text representations in the speech head: one second of speech corresponds to hundreds of tokens, far exceeding the amount of text information humans typically express per second. In LLMs, multi-token prediction (MTP) (Gloeckle et al. 2024; Cai et al. 2024; DeepSeek-AI et al. 2024) is often used to accelerate inference or enrich hidden representations by enabling a single hidden state to predict multiple future tokens. To balance the information density between speech and text, we introduce MTP into SLMs, allowing each speech hidden vector to decode multiple speech tokens (for both coupled and decoupled tokenizers). We experimented with various compression ratios, ranging from one token per hidden state to twelve tokens per state. The MTP not only improved speech decoding speed by up to 12 but also significantly enhanced speech synthesis quality (the Word Error Rate decreased from 6.07 to 3.01).
In addition, we propose a speaker-aware speech generation paradigm, introducing speaker identity features into the speech-language context to guide and control the speaker timbre in synthesized speech. To comprehensively evaluate SLMs’ understanding and generalization of the general knowledge captured by LLMs, as well as their speaker consistency, we construct a role-playing knowledge QA task. Specifically, we use CosyVoice2 (2024b) to convert answers from a text-based knowledge QA dataset—TriviaQA (2017) into speech, assigning each answer to one of 15 Genshin characters. The resulting RoleTriviaQA dataset contains 138K/500/2,426 samples for training/validation/test, respectively. We reserve five characters and the WebQuestions (2025) test set for out-of-domain evaluation of both speaker consistency and knowledge QA performance (i.e., evaluating generalization). Results show that the decoupled architecture enables cooperative improvement in knowledge QA tasks and speaker timbre, and significantly outperforms other baselines.
@@ 2 Speech-Language Model
A typical speech-language model (SLM) is usually an extension of a successful large language model (LLM) capable of understanding and generating both text and speech signals. It relies on a speech-language tokenizer to convert raw text or audio into discrete tokens, a decoder-only Transformer to learn cross-modal context and alignment, a language head to predict text, and a speech head to predict speech tokens then synthesize continuous audio signals.
@@ 2.1 Core Components
has achieved great success as a foundation model for text understanding and generation. After unifying the tokenization of text and speech, it can further learn joint speech-text context and perform cross-modal alignment. Its parameters are typically initialized from a well-pretrained LLM and then adapted to multimodal contexts. Given a sequence of input tokens , it produces a hidden representation vector for each token,
@@
A parameter matrix maps the hidden vector to the vocabulary space of size . During training, the model is optimized using the cross-entropy loss: , and during inference, the next token is predicted by either taking the or sampling.
Speech head follows a similar formal definition, training objective, and inference process as the language head,
@@ 2.2 Multi Token Prediction
We observe that text representations are significantly more information-dense than speech. Typically, one second of speech is decoded into several hundred tokens, while it rarely conveys more than ~20 text tokens’ worth of information. This asymmetry in information density can make cross-modal alignment difficult to learn. On the other hand, speech tokens with excessively low information density can also hinder SLM inference efficiency, as generating one second of speech may take significantly longer than one second. Therefore, balancing the information density between speech and text representations becomes a critical challenge.
Multi-token prediction (MTP) (Gloeckle et al. 2024; Cai et al. 2024; Liu et al. 2024) is commonly used in LLMs to accelerate inference or enrich hidden representations (e.g., by requiring to decode not only but also future tokens like ). To compress the information density of speech representations, we introduce MTP into the SLM framework. As depicted in Figure 2, compared with Next Token Prediction (NTP), MTP enables each hidden vector to decode multiple speech tokens, whether coupled or decoupled, thus alleviates the imbalance in information density between speech and text.
@@
The loss / is also extended to the average over all tokens within the group,
@@
where the is the concatenate function.
@@ 2.3 Speaker-Aware Speech Generation
Model |
Nq |
UTMOS | WER | SIM | |||||||
Seen | Unseen | |||||||||||||
Coupled Tokenizers trained with SLM | ||||||||||||||
Encodec (2023) | 1024 | 4 | 300 | 0.38 | 2.86 | 17.00 | 0.18 | 0.10 | ||||||
Encodec (2023) | 1024 | 8 | 600 | 0.24 | 3.50 | 8.14 | 0.18 | 0.10 | ||||||
WavTokenizer (2025) | 4096 | 1 | 75 | 0.38 | 2.80 | 89.60 | 0.06 | 0.04 | ||||||
WavTokenizer-v2 (2025) | 4096 | 1 | 75 | 0.65 | 3.16 | 49.24 | 0.11 | 0.08 | ||||||
StableCodec (2024) | 15625 | 2 | 50 | 0.98 | 3.90 | 13.04 | 0.17 | 0.14 | ||||||
BigCodec (2024) | 8192 | 1 | 80 | 0.70 | 3.93 | 13.63 | 0.18 | 0.10 | ||||||
Semi-Decoupled Tokenizers trained with SLM | ||||||||||||||
SpeechTokenizer (2024) | 1024 | 5 | 250 | 0.99 | 3.86 | 13.13 | 0.16 | 0.13 | ||||||
SpeechTokenizer (2024) | 1024 | 8 | 400 | 0.90 | 4.05 | 9.95 | 0.14 | 0.14 | ||||||
Decoupled Tokenizers trained with SLM | ||||||||||||||
FACodec-NTP-1H | 1024 | 3 | 240 | 0.87 | 3.93 | 6.07 | 0.50 | 0.49 | ||||||
FACodec-MTP-3H | 1024 | 3 | 80 | 0.82 | 3.83 | 5.84 | 0.47 | 0.47 | ||||||
FACodec-MTP-6H | 1024 | 3 | 40 | 0.92 | 3.75 | 4.36 | 0.47 | 0.47 | ||||||
@@ 3 Role-Playing Knowledge QA Task
However, a critical gap exists: no open-source dataset assesses both correctness and timbre similarity in speech knowledge QA. Other models featuring speech QA (Défossez et al. 2024) trained and evaluated on a large-scale data which has not been released yet. To bridge this gap and stimulate SLMs’ intrinsic knowledge with substantial training (Zhang et al. 2023), we propose RoleTriviaQA, an open-source dataset for benchmarking SLMs on role embodiment and knowledgeable spoken responses. Similar with (Zhang et al. 2023), we adopt the data format (see Figure 3), which enables essential text-speech knowledge alignment for high-quality role-playing.
@@ 3.1 Data Construction
We created the RoleTriviaQA dataset by selecting 15 distinct voices from Genshin Impact555https://genshin.hoyoverse.com/en roles (10 seen: 5 Male/5 Female for training and evaluation; 5 unseen: 2 Male/3 Female for evaluation only). Then we synthesize speech for TriviaQA (Joshi et al. 2017) QA-pairs using CosyVoice 2 (Du et al. 2024b) after filtering high-quality reference speeches. This pipeline yields 138,384 training samples and a test set with 1,500 in-domain (TriviaQA) and 926 out-of-domain (OpenAudioBench (Li et al. 2025)) samples.Further details of RoleTriviaQA are in Appendix H.
@@ 3.2 Training Strategy
To improve cross-modal alignment, we conduct a large-scale dataset Emilia-3.5. First, we filter Emilia (2024) (a large-scale, real-world speech collection with diverse emotions and styles) for samples scoring at least 3.5 on DNSMOS (2022). Next, we merge the standard LibriTTS (2019) training set. The resulting corpus comprises 2,495 hours of high-quality speech. Moreover, To enable bidirectional alignment of text and speech modalities, we conduct pretrain on both TTS and ASR tasks. Specifically, we adopt as the TTS data format and for ASR. The pretaining loss is formulated as:
where the predicted token could be a single token or a multi-token group (i.e., ).
@@
The SFT loss is defined as follows:
where the loss is calculated across both the text answer and the speech answer.
@@ 4 Experiment
Models |
In-Domain (RoleTriviaQA) | Out-of-Domain (Web Questions) | ||||||||||||
EM(%) | F1(%) | SIM | EM | F1 | SIM | |||||||||||
Seen | Unseen | All | Seen | Unseen | All | |||||||||||
Qwen2.5-0.5B-Instruct | LLM | - | 7.1 | 17.3 | - | - | - | 1.1 | 14.0 | - | - | - | ||||
WavTokenizer-v2 | SLM | Coupled | 5.1 | 12.5 | 0.22 | 0.18 | 0.20 | 1.1 | 3.9 | 0.21 | 0.18 | 0.20 | ||||
BigCodec | SLM | Coupled | 2.9 | 9.4 | 0.23 | 0.17 | 0.21 | 1.0 | 4.2 | 0.23 | 0.17 | 0.21 | ||||
StableCodec | SLM | Coupled | 7.5 | 13.8 | 0.18 | 0.14 | 0.17 | 0.6 | 3.6 | 0.17 | 0.14 | 0.16 | ||||
SpeechTokenizer | SLM | Semi-Decoupled | 3.4 | 9.1 | 0.19 | 0.14 | 0.17 | 0.3 | 2.6 | 0.20 | 0.14 | 0.18 | ||||
@@ 4.1 Effect of Speech Tokenizers
First, coupled architectures exhibit severe convergence challenges during SLM training (shown in Fig. 4), resulting in generally low success rates of speech synthesis (70%). The notable exception is StableCodec, whose large codebook (15,625 entries) and low token rate (50 TPS) establish dual compensation mechanisms that mitigate inherent limitations of coupled architectures. Second, the semi-decoupled and decoupled architectures show clear advantages. The semi-decoupled architecture (SpeechTokenizer) achieves the best speech quality (UTMOS), while the decoupled architecture works optimally in WER and SIM. Third, the decoupled architectures’ SIM between the training and test sets is consistently larger and significantly better than the other baselines, suggesting superior generalization capabilities, especially for speaker timbre.
Overall, decoupled tokenizer and SLM joint modeling is better, and has better controllability and scalability.
@@
As shown in Table 1, to investigate the effectiveness of MTP under decoupled architecture, we also conduct experiments comparing models trained with different compression ratios.
First, the MTP-3H model is competitive with the NTP-1H, even with compression. Moreover, it has also surpassed all of the semi-decoupled and coupled baselines in WER and SIM. Second, when increasing compression from (MTP-3H) to (MTP-12H), WER improves by 48% (5.84 to 3.01), with stable SIM and UTMOS experiencing a slight decline. Notably, under 12 compression, our model achieves speech synthesis with a 100% success rate, and the WER (3.01) is comparable to the ground truth. These results reveal that speech length compression based on the decoupled tokenizer effectively mitigates the frequency mismatch between text (~20Hz) and speech (240Hz) modalities, enhancing cross-modal alignment performance.

### v3

@@ Abstract
Speech-language models (SLMs) offer a promising path toward unifying speech and text understanding and generation. However, challenges remain in achieving effective cross-modal alignment and high-quality speech generation. In this work, we systematically investigate the role of speech tokenizer designs in LLM-centric SLMs, augmented by speech heads and speaker modeling. We compare coupled, semi-decoupled, and fully decoupled speech tokenizers under a fair SLM framework and find that decoupled tokenization significantly improves alignment and synthesis quality. To address the information density mismatch between speech and text, we introduce multi-token prediction (MTP) into SLMs, enabling each hidden state to decode multiple speech tokens. This results in up to faster decoding and a substantial reduction in word error rate (from 6.07 to 3.01). Furthermore, we propose a speaker-aware generation paradigm and introduce RoleTriviaQA, a large-scale role-playing knowledge QA benchmark with diverse speaker identities. Experiments demonstrate that our methods enhance both knowledge understanding and speaker consistency.
Code — https://github.com/cnxupupup/SLM-Decoupled-MTP
Demo — https://cnxupupup.github.io/SLM-Decoupled-MTP-Demo
Dataset — https://huggingface.co/datasets/cnxup/RoleTriviaQA
@@ 1 Introduction
In recent years, large language models (LLMs) have achieved remarkable progress in text understanding and generation (Grattafiori et al. 2024; Qwen 2025), fueling rapid advancements in multimodal models such as vision-language models (Liu et al. 2023), vision-language-action models (Kim et al. 2024), and speech-language models (SLMs) (Chu et al. 2023; Li et al. 2025). Among them, SLMs are particularly valuable for applications like natural and fluent human-computer dialogue (Défossez et al. 2024; Xu et al. 2025) and personalized speech generation (Du et al. 2024a, b). The core challenge lies in generalizing the language understanding and generation capabilities of LLMs to incorporate and integrate speech understanding and generation.
Current approaches for connecting text and speech in SLMs can be broadly categorized into two perspectives. Speech-centric models (Du et al. 2024a, b; Anastassiou et al. 2024) are built upon mature speech generation architectures, with LLMs introduced as a conditional context. However, while this approach achieves high-quality speech synthesis, it struggles to fully leverage LLMs’ rich world knowledge and powerful capabilities. LLM-centric models (Défossez et al. 2024; Zhang et al. 2023; Zhan et al. 2024) take LLMs as the backbone and extend them with speech interaction. In theory, it fully utilizes LLMs’ strengths (e.g., intent understanding, instruction following, planning, and decision-making). However, the quality of speech generation often suffers due to challenges in cross-modal alignment. This work focuses on the LLM-centric SLMs, aiming to explore how core components affect cross-modal alignment and the quality of generated speech.
@@
We observe a mismatch in information density between speech and text representations in the speech head: one second of speech corresponds to hundreds of tokens, far exceeding the amount of text information humans typically express per second. In LLMs, multi-token prediction (MTP) (Gloeckle et al. 2024; Cai et al. 2024; DeepSeek-AI 2025) is often used to accelerate inference or enrich hidden representations by enabling a single hidden state to predict multiple future tokens. To balance the information density between speech and text, we introduce MTP into SLMs, allowing each speech hidden vector to decode multiple speech tokens (for both coupled and decoupled tokenizers). We experimented with various compression ratios, ranging from one token per hidden state to twelve tokens per state. The MTP not only improved speech decoding speed by up to 12 but also significantly enhanced speech synthesis quality (the Word Error Rate decreased from 6.07 to 3.01).
In addition, we propose a speaker-aware speech generation paradigm that introduces speaker identity features into the speech-language context to guide and control the speaker’s timbre in synthesized speech. To comprehensively evaluate SLMs’ understanding and generalization of the general knowledge captured by LLMs, as well as their speaker consistency, we construct a role-playing knowledge QA task. Specifically, we use CosyVoice2 (2024b) to convert answers from a text-based knowledge QA dataset—TriviaQA (2017) into speech, assigning each answer to one of 15 Genshin characters. The resulting RoleTriviaQA dataset contains 138K/0.3K/2.4K samples for training/validation/test, respectively. We reserve five characters and the WebQuestions (2025) test set for out-of-domain evaluation of both speaker consistency and knowledge QA performance (i.e., evaluating generalization). Results show that the decoupled architecture enables cooperative improvement in knowledge QA tasks and speaker timbre, and significantly outperforms other baselines.
@@ 2 Speech-Language Model
A typical speech-language model (SLM) is usually an extension of a successful large language model (LLM) capable of understanding and generating both text and speech signals. It relies on a speech-language tokenizer to convert raw text or audio into discrete tokens, a decoder-only Transformer to learn cross-modal context and alignment, a language head to predict text, and a speech head to predict speech tokens and then synthesize continuous audio signals.
@@ 2.1 Core Components
has achieved great success as a foundation model for text understanding and generation. After unifying the tokenization of text and speech, it can further learn joint speech-text context and perform cross-modal alignment. Its parameters are typically initialized from a well-pretrained LLM and then adapted to multimodal contexts. Given a sequence of input tokens , it produces a hidden representation vector and for each token,
@@
A parameter matrix maps the hidden vector to the vocabulary space of size . During training, the model is optimized using the cross-entropy loss: , and during inference, the next token is predicted by either taking the or sampling. Speech head follows a similar formal definition, training objective, and inference process as the language head,
@@ 2.2 Multi Token Prediction
We observe that text representations are significantly more information-dense than speech. Typically, one second of speech is decoded into several hundred tokens, while it rarely conveys more than 20 text tokens’ worth of information. This asymmetry in information density can make cross-modal alignment difficult to learn. On the other hand, speech tokens with excessively low information density can also hinder SLM inference efficiency, as generating one second of speech may take significantly longer than one second. Therefore, balancing the information density between speech and text representations becomes a critical challenge.
Multi-token prediction (MTP) (Gloeckle et al. 2024; Cai et al. 2024; DeepSeek-AI 2025) is commonly used in LLMs to accelerate inference or enrich hidden representations (e.g., by requiring to decode not only but also future tokens like ). To compress the information density of speech representations, we introduce MTP into the SLM framework. As depicted in Figure 2, compared with Next Token Prediction (NTP), MTP enables each hidden vector to decode multiple speech tokens, whether coupled or decoupled, thus alleviating the imbalance in information density between speech and text.
@@
The loss / is also extended to the average over all tokens within the group:
@@
where the is the concatenation function.
@@ 2.3 Speaker-Aware Speech Generation
| Model | Codebook Size | Nq | Token Rate (TPS) | Success Rate (SR) | UTMOS | WER | SIM | |
| Seen | Unseen | |||||||
| Coupled Tokenizers trained with SLM | ||||||||
| Encodec (2023) | 1024 | 4 | 300 | 0.38 | 2.86 | 17.00 | 0.18 | 0.10 |
| Encodec (2023) | 1024 | 8 | 600 | 0.24 | 3.50 | 8.14 | 0.18 | 0.10 |
| WavTokenizer (2025) | 4096 | 1 | 75 | 0.38 | 2.80 | 89.60 | 0.06 | 0.04 |
| WavTokenizer-v2 (2025) | 4096 | 1 | 75 | 0.65 | 3.16 | 49.24 | 0.11 | 0.08 |
| StableCodec (2025) | 15625 | 2 | 50 | 0.98 | 3.90 | 13.04 | 0.17 | 0.14 |
| BigCodec (2024) | 8192 | 1 | 80 | 0.70 | 3.93 | 13.63 | 0.18 | 0.10 |
| Semi-Decoupled Tokenizers trained with SLM | ||||||||
| SpeechTokenizer (2024) | 1024 | 5 | 250 | 0.99 | 3.86 | 13.13 | 0.16 | 0.13 |
| SpeechTokenizer (2024) | 1024 | 8 | 400 | 0.90 | 4.05 | 9.95 | 0.14 | 0.14 |
| Decoupled Tokenizers trained with SLM | ||||||||
| FACodec-NTP-1H | 1024 | 3 | 240 | 0.87 | 3.93 | 6.07 | 0.50 | 0.49 |
| FACodec-MTP-3H | 1024 | 3 | 80 | 0.82 | 3.83 | 5.84 | 0.47 | 0.47 |
| FACodec-MTP-6H | 1024 | 3 | 40 | 0.92 | 3.75 | 4.36 | 0.47 | 0.47 |
@@ 3 Role-Playing Knowledge QA Task
However, a critical gap exists: no open-source dataset assesses both correctness and timbre similarity in speech knowledge QA. Other models featuring speech QA (Défossez et al. 2024) were trained and evaluated on a large-scale dataset, which has not been released yet.
To bridge this gap and stimulate SLMs’ intrinsic knowledge with substantial training (Zhang et al. 2023), we propose RoleTriviaQA, an open-source dataset for benchmarking SLMs on role embodiment and knowledgeable spoken responses. Similar with (Zhang et al. 2023), we adopt the data format (see Figure 3), which enables essential text-speech knowledge alignment for high-quality role-playing.
@@ 3.1 Data Construction
We created the RoleTriviaQA dataset by selecting 15 distinct voices from Genshin Impact roles (10 seen: 5 Male/5 Female for training and evaluation; 5 unseen: 2 Male/3 Female for evaluation only). Then we synthesize speech for TriviaQA (Joshi et al. 2017) QA-pairs using CosyVoice 2 (Du et al. 2024b) after filtering high-quality reference speeches. This pipeline yields 138,384 training samples and a test set with 1,500 in-domain (ID, from TriviaQA) and 926 out-of-domain (OOD, from OpenAudioBench (Li et al. 2025)) samples. Further details of RoleTriviaQA are in Appendix H.
@@ 3.2 Training Strategy
To improve cross-modal alignment, we conduct a large-scale dataset Emilia-3.5. First, we filter Emilia (2024) (a large-scale, real-world speech collection with diverse emotions and styles) for samples scoring at least 3.5 on DNSMOS (2022). Next, we merge the standard LibriTTS (2019) training set. The resulting corpus comprises 2,495 hours of high-quality speech. Moreover, to enable bidirectional alignment of text and speech modalities, we conduct pre-training on both TTS and ASR tasks. Specifically, we adopt as the TTS data format and for ASR. The pretaining loss is formulated as: where the predicted token could be a single token or a multi-token group (i.e., ).
@@
The SFT loss is defined as follows: where the loss is calculated across both the text answer and the speech answer.
@@ 4 Experiment
| Models | Model Type | Tokenizer Type | In-Domain (RoleTriviaQA) | Out-of-Domain (Web Questions) | ||||||||
| EM(%) | F1(%) | SIM | EM | F1 | SIM | |||||||
| Seen | Unseen | All | Seen | Unseen | All | |||||||
| Qwen2.5-0.5B-Instruct | LLM | - | 7.1 | 17.3 | - | - | - | 1.1 | 14.0 | - | - | - |
| WavTokenizer-v2 | SLM | Coupled | 5.1 | 12.5 | 0.22 | 0.18 | 0.20 | 1.1 | 3.9 | 0.21 | 0.18 | 0.20 |
| BigCodec | SLM | Coupled | 2.9 | 9.4 | 0.23 | 0.17 | 0.21 | 1.0 | 4.2 | 0.23 | 0.17 | 0.21 |
| StableCodec | SLM | Coupled | 7.5 | 13.8 | 0.18 | 0.14 | 0.17 | 0.6 | 3.6 | 0.17 | 0.14 | 0.16 |
| SpeechTokenizer | SLM | Semi-Decoupled | 3.4 | 9.1 | 0.19 | 0.14 | 0.17 | 0.3 | 2.6 | 0.20 | 0.14 | 0.18 |
@@ 4.1 Effect of Speech Tokenizers
First, coupled architectures exhibit severe convergence challenges during SLM training (see Appendix C), resulting in generally low success rates of speech synthesis (70%). The notable exception is StableCodec, whose large codebook and low token rate establish dual compensation mechanisms that mitigate inherent limitations of coupled architectures. Second, the semi-decoupled and decoupled architectures show clear advantages. The semi-decoupled architecture achieves the best speech quality, while the decoupled architecture works optimally in WER and SIM. Third, the decoupled architectures’ SIM between the training and test sets is consistently larger and significantly better than the other baselines, suggesting superior generalization capabilities, especially for speaker timbre.
Overall, decoupled tokenizer and SLM joint modeling is better and has better controllability and scalability.
@@
As illustrated in Table 1, to investigate the effectiveness of MTP under decoupled architecture, we also conduct experiments comparing models trained with different compression ratios.
First, the MTP-3H model is competitive with the NTP one, even with compression. Moreover, it has also surpassed all of the semi-decoupled and coupled baselines in WER and SIM. Second, when increasing compression from (MTP-3H) to (MTP-12H), WER improves by 48% (5.84 to 3.01), with stable SIM and UTMOS experiencing a slight decline. Notably, under compression, our model achieves speech synthesis with a 100% success rate, and the WER (3.01) is comparable to the ground truth. These results reveal that speech length compression based on the decoupled tokenizer effectively mitigates the frequency mismatch between text (20Hz) and speech (240Hz) modalities, enhancing cross-modal alignment performance.